A 100-W lightbulb has a resistance of about 12 Ω when cold (20 ∘C) and 128 Ω when on (hot).

Estimate the temperature of the filament when hot assuming an average temperature coefficient of resistivity α=0.0045(C)−1.
Answer #1

R = R° [ 1 + α ( T - T°)]
R° = 12 Ω
R = 128 Ω
T° = 20°
α = 4.5 10^-3
128 = 12 [ 1 + 4.5 10-3 ( T - 20)]
128 = 12 + 54 10^-3 ( T -20)
128 - 12 = 54 10^-3 T - 1.08
117.08 = 54 10^-3 T ---- T = 117.08 / 54 10^-3
T = 117.08 * 10^3 / 54 = 2168.14. °C